Can someone help me with this problem?
The following table gives the winning percentages of the tow 5 coaches inNFL history. What fraction of his games has each coach won?
Don Shula - 0.676
George Halas - 0.672
Tom Landry - 0.606
Curly Lambeau - 0.624
Paul Brown - 0.667

Answer by stanbon(75887)

What fraction of his games has each coach won?
Don Shula - 0.676= 676/1000
George Halas - 0.672=672/1000
Tom Landry - 0.606=606/1000
Curly Lambeau - 0.624=624/1000
Paul Brown - 0.667=667/1000
Some of these you can further reduce, but each
one does answer your question.
Cheers,
Stan H.